Given the following data, what formulas would one use to complete a spreadsheet that shows the related CVP Analysis computations? Thanks!

Course Documents-MA \ D Study Guide for Test-2-, ?D Chapter 5 CVP Excel x \. x ox?0.32041458693813785 1527375348486 Caibri Blu- .2.?. Alignment Number Cordboral Formetas. Cell Cells Editing Formatting TableStyles Styies - A1 ¥ | i | X ? | Given the following information complete a CVP analysis 1 Given the following information complete a CVlanalysis 2 for JPL, Inc 4 Unit sales 11,200 units Selling price per unit 6 Variable expenses per unit $75 per unit 45 per unit Fixed expenses $210,000 9 Use the data to answer the following. 11 1. Compute the CM ratio and variable expense ratio 2 Selling price per unit 1 Variable expenses per unit 14 Contribution margin per unit per unit per unit per unit 16 CM ratio 17 Varlable expense ratio 19 2. Compute the break-even paint 20 Break-even in unit sales 21 Break-even in dollar sales units 23 3.Compute the margin of s n doll Sheeti eBook & Resources
Answer #1

Answer

  • All working forms part of the answer
  • Please follow the ‘working equation column’ to understand computation of figures

A

Unit Sales

11200

units

B

Selling Price per unit

$                                   75.00

per unit

C

Variable expenses per unit

$                                   45.00

per unit

D

Fixed Expenses

$                       2,10,000.00

B

Selling Price per unit

$                                   75.00

per unit

C

Variable expenses per unit

$                                   45.00

per unit

E= B - C

Contribution margin per unit

$                                   30.00

per unit

F = E/B

CM ratio

40%

G = C/B

Variable Expense ratio

60%

H = D/E

Break Even in Unit Sales

7000

I = D/F

Break Even in Dollar Sales

$                       5,25,000.00

J = 11200 x $75

Total Sales

$                       8,40,000.00

K = I

Break Even in Dollar Sales

$                       5,25,000.00

L = J - K

Margin of Safety in dollars

$                       3,15,000.00

M = J

Sales

$                       8,40,000.00

N = 11200 x $45

Variable expenses

$                       5,04,000.00

O = M - N

Contribution margin

$                       3,36,000.00

P = D

Fixed Expenses

$                       2,10,000.00

Q = O - P

Net Operating Income

$                       1,26,000.00

R = O/Q

Degree of operating Leverage

2.67